OpenStudy (anonymous):

solve the system by the substitution method x-7y=8 2x-14y=11

OpenStudy (anonymous):

x-7y=8 Rearranging above we get x = 8 + 7y Substituting this value of x in second equation 2(8 + 7y) -14y=11 16 +14y - 14y = 11 This gives 16 = 11 which is wrong Therefore the two given equations do not have a solution.